OpenStudy (kkbrookly):

Will give medal and fan (:

OpenStudy (kkbrookly):

@freckles @Michele_Laino @Nnesha

OpenStudy (freckles):

hint sin(x) and csc(x) are reciprocals of one another and cos(x) and sec(x) are reciprocals of one another and tan(x) and cot(x) are reciprocals of one another

OpenStudy (freckles):

u and 1/u are reciprocals and their product is 1 when u is not 0 of course

OpenStudy (freckles):

\[u \cdot \frac{1}{u}=\frac{u}{u}=1 , u \neq 0\]

OpenStudy (kkbrookly):

Meaning I replace them with their reciprocal or what do I do with it?

OpenStudy (freckles):

the product of reciprocals is 1 on their domain

OpenStudy (freckles):

sin(x)*csc(x)=? cos(x)*sec(x)=? tan(x)*cot(x)=? assuming of course x is in the domain

OpenStudy (freckles):

you do know that csc(x)=1/sin(x) and cos(x)=1/sec(x) and tan(x)=1/cot(x)?

OpenStudy (anonymous):

whwt is "tan" and "cos"?

OpenStudy (freckles):

\[\csc(x)=\frac{1}{\sin(x)} \text{ implies } \sin(x) \cdot \csc(x)=1 \] and so on

OpenStudy (kkbrookly):

So all of those equal one then?

OpenStudy (freckles):

yes the product of reciprocals is 1

OpenStudy (freckles):

so you can reduce your question to what is the following simplify to: 1(1)(1)=?

OpenStudy (kkbrookly):

So, the answer is just 1?

OpenStudy (freckles):

yep it is 1 assuming x is an element of our domain of our original expression

OpenStudy (kkbrookly):

She never stated what x was so I'm guessing that's it. Are you able to help me with one more that's dividing?

OpenStudy (freckles):

sure usually we are under the assumption that the variable we are using is in the domain of the expression

OpenStudy (freckles):

it x is not an element of the domain the output would have been undefined

OpenStudy (freckles):

but anyways yes sure what is the second question

OpenStudy (freckles):

\[\sec^2(x)-\frac{\tan^2(x)}{\csc(x)}?\]

OpenStudy (freckles):

or \[\frac{\sec^2(x)-\tan^2(x)}{\csc(x)}\]?

OpenStudy (kkbrookly):

The second one

OpenStudy (freckles):

do you know your Pythagorean identities?

OpenStudy (freckles):

\[\sin^2(x)+\cos^2(x)=1 \\ 1+\cot^2(x)=\csc^2(x) \\ \tan^2(x)+1=\sec^2(x)\]

OpenStudy (freckles):

do these look familiar ?

OpenStudy (kkbrookly):

Yes, they look familiar.

OpenStudy (freckles):

tried to apply that last equation I wrote

OpenStudy (kkbrookly):

So do I do tan^2(x)+1=sec^2(x)/csc x or how do I even figure out the csc x with this equation?

OpenStudy (freckles):

I'm not sure how you applied the equation I asked you to apply. What is your numerator simplify to upon using the equation I asked you to use?

OpenStudy (kkbrookly):

The numerator would be 1, right? Because the tan^2x would cancel out

OpenStudy (freckles):

that is exactly right so you would have 1/csc(x)

OpenStudy (freckles):

and I actually gave you what 1/csc(x) equals to earlier

OpenStudy (freckles):

recall that sin(x) and csc(x) are again reciprocals of one another

OpenStudy (kkbrookly):

You said that csc(x)=1/sin(x) so would the answer be sin(x)?

OpenStudy (freckles):

yeah yeah! :)

OpenStudy (freckles):

\[\frac{1}{\csc(x)}=\sin(x) \text{ because } \frac{1}{\sin(x)}=\csc(x)\]

OpenStudy (kkbrookly):

Thank you so much! You explained it better than any other person on here

OpenStudy (freckles):

i tried and I bet there are some people who can explain it better than me :p

OpenStudy (freckles):

but thanks :)

OpenStudy (kkbrookly):

But you were the person that actually took the time to help so I'm really grateful. Thanks again (: